The heat of melting for Cu is 13,260 J/mol and the equilibrium melting temperature of Cu is 1358 K. Assuming the constant pressure heat capacities of solid Cu and liquid Cu are the same, 24.0 J/(mol K) What is the entropy change for the surroundings, which is also at 1000 K, during solidification from one mole of liquid Cu to one mole solid Cu in the system at 1000 K?
Answer #1

There are two processes involved here , the entropy change can be given as :-

\Delta S =\Delta H/T_{m} + C_{p}ln[T_{2}/T_{1}]

Delta H = 13260 J/mol

melting temperature = 1358 K

initial temperature =1358 K

Final temperature = 1000 K

Cp = 24..0 J/mol putting these values

\Delta S =13260/1358 + 24\times ln[1000/1358]

\Delta S =9.764 -7.344 = 2.419 J/K mol